How many liters of water should a chemist mix with 10 liters of a 24% sodium solution to get an 8% sodium solution?
So to line it up, I have:
All L liters.
10*24 + L*0 = (10+L)*10

10L + 24 = 240
10l = 264
26.4... Im doing something wrong and I dont know what.
Please help me. Thank you!

Answer by Alan3354(69443)

How many liters of water should a chemist mix with 10 liters of a 24% sodium solution to get an 8% sodium solution?
So to line it up, I have:
All L liters.
10*24 + L*0 = (10+L)*10 ******** You said 8%
10*24 + L*0 = (10+L)*8
240 = 80 + 8L
L = 20 liters
=================================

10L + 24 = 240 ??????
10l = 264 ????????
26.4... Im doing something wrong and I dont know what.
You did most things wrong.
-------------------------
Please help me. Thank you!